# Generalized Frobenius Manifold Structures on the Orbit Spaces of Affine Weyl Groups I

### v1

@@ 4.2. The polynomiality property
In other words, upon taking coordinates $(t^{1},\cdots,t^{\ell})$ on $\overline{\mathcal{M}}$ and $(z^{1},\cdots,z^{\ell})$ on $\mathcal{M}\setminus D$, the projection $\overline{\pi}$ coincides with the natural map $\mathrm{Spec}(\mathbb{C}[t^{1},\cdots,t^{\ell}])\to\mathrm{Spec}(\mathbb{C}[z^{1},\cdots,z^{\ell}])$.
4.3. The Monodromy group of $\mathcal{M}(R,\omega)$
The definition of the monodromy group of a Frobenius manifold was given by Dubrovin in [9], and it can be extended directly to generalized Frobenius manifolds. Let us first recall this defintion.
Let $M$ be an $\ell$-dimensional generalized Frobenius manifold, of which the flat metric is $\eta$ and the intersection form is $g$, and let $t^{1},\cdots,t^{\ell}$ be flat coordinates of $\eta$. Consider the system of equations
$\frac{\partial^{2}u}{\partial t^{\alpha}\partial t^{\beta}}=\Gamma_{\alpha\beta}^{\epsilon}\frac{\partial u}{\partial t^{\epsilon}},\quad\alpha,\beta=1,\dots,\ell$ (4.26)
for the flat coordinates of $g$. Here $\Gamma_{\alpha\beta}^{\epsilon}$ are the Christoffel symbols of the Levi-Civita connection of $g$. Denote $\Sigma=\{t\in M~{}|~{}\det g(t)=0\}$, then the solutions $u$ live on the universal covering $\widehat{M\setminus\Sigma}$ of $M\setminus\Sigma$. Define the solution space
[…]
then we know that $\dim\mathcal{S}=n$, and there exist flat coordinates $u^{1},\cdots,u^{\ell}$ of $g$ such that
$\mathcal{S}=\mathrm{span}\{u^{1},\cdots,u^{\ell}\}.$ (4.28)
For any loop $\gamma\in\pi_{1}(M\setminus\Sigma)$, we have a transformation
$\gamma^{*}u^{\alpha}=A_{\beta}^{\alpha}(\gamma)u^{\beta}+B^{\alpha}(\gamma),\quad\alpha=1,\dots,\ell$ (4.29)
on $\mathcal{S}$ by analytic continuation. Denote by $E^{\ell}$ the dual space of $\mathcal{S}$, then it has coordinates $u^{1},\cdots,u^{\ell}$. Hence the transformation (4.29) yields a transformation on $E^{\ell}$, which we also denote by $\gamma^{*}$. Since $E^{\ell}$ is an $\ell$-dimensional complexified Euclidean space $\mathbb{R}^{\ell}\otimes\mathbb{C}$ equipped with the metric $g$, and the action by $\pi_{1}(M\setminus\Sigma)$ preserves the metric $g$, this transformation on $E^{\ell}$ is an isometry. Hence we have obtained a representation
$r\colon\pi_{1}(M\setminus\Sigma)\to\mathrm{Iso}(E^{\ell})\cong O(\ell)\rtimes\mathbb{C}^{\ell},\quad\gamma\mapsto\gamma^{*}.$ (4.30)
@@ 4.3. The Monodromy group of $\mathcal{M}(R,\omega)$
$\mathrm{Mono}(M)=r(\pi_{1}(\mathcal{M}\setminus\Sigma))\leq O(\ell)\rtimes\mathbb{C}^{\ell}.$ (4.31)
Now we proceed to compute the monodromy group of the generalized Frobenius manifold ¯$\overline{\mathcal{M}}=\mathcal{M}(R,\omega)$. Let
$\overline{\Sigma}=\{t\in\overline{\mathcal{M}}\mid\det\left(g^{\alpha\beta}(t)\right)=0\},$
and $\mathcal{U}(\mathcal{M})$ be the universal covering space of ¯$\overline{\mathcal{M}}\setminus\overline{\Sigma}$. The flat coordiantes of $g$ are $x^{1}\bigr{|}_{\lambda=0},\cdots,x^{\ell}\bigr{|}_{\lambda=0}$, and they are functions on $\mathcal{U}(\mathcal{M})$. Moreover, $x^{1}\bigr{|}_{\lambda=0},\cdots,x^{\ell}\bigr{|}_{\lambda=0}$ are also coordinates of $V\otimes\mathbb{C}$, so we can view the complexified Euclidean space $E^{\ell}$ as $V\otimes\mathbb{C}$. Hence we obtain a representation
$r\colon\pi_{1}(\overline{\mathcal{M}}\setminus\overline{\Sigma})\to\mathrm{Iso}(V\otimes\mathbb{C})\cong O(\ell)\rtimes\mathbb{C}^{\ell},\quad\gamma\mapsto\gamma^{*},$ (4.32)
@@
Here the image $r(\pi_{1}(\overline{\mathcal{M}}\setminus\overline{\Sigma}))$ is the monodromy group of $\overline{\mathcal{M}}$. However, this group is difficult to compute, and we will instead consider a group bigger than it.
@@
$\overline{\pi}\colon\overline{\mathcal{M}}\to\mathcal{M}\setminus D$ (4.34)
is nondegerated, so from
$\det\left(g^{\alpha\beta}(t)\right)=\det\left(g^{ij}(z)\right)\det\left(\frac{\partial t^{\alpha}}{\partial z^{j}}\right)$ (4.35)
we know that $\left(g^{\alpha\beta}(t)\right)$ degenerates if and only if $\left(g^{ij}(z(t)\right)$ degenerates. In other words, denote
$\Sigma^{\prime}=\{z\in\mathcal{M}\setminus D\mid\det\left(g^{ij}(z)\right)=0\},$
then we have $(\overline{\pi})^{-1}(\Sigma^{\prime})=\overline{\Sigma}$, and the map
$\overline{\pi}:\overline{\mathcal{M}}\setminus\overline{\Sigma}\to\mathcal{M}\setminus(D\cup\Sigma^{\prime})$ (4.36)
is also a covering map, hence
$\pi_{1}(\overline{\mathcal{M}}\setminus\overline{\Sigma})\leq\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})).$ (4.37)
Since ¯$\overline{\mathcal{M}}\setminus\overline{\Sigma}$ is a covering space of $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$, they share the same universal covering space $\mathcal{U}(\mathcal{M})$, and both $\pi_{1}(\overline{\mathcal{M}}\setminus\overline{\Sigma})$ and $\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))$ act on $\mathcal{U}(\mathcal{M})$. Note that $g^{ij}(z)$ are polynomials of $z^{1},\cdots,z^{\ell}$, so in particular they are single valued functions on $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$. This implies that the action of $\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))$ on $\mathcal{U}(\mathcal{M})$ also leaves the metric $g$ invariant, so the representation (4.32) can be extended to a representation
$r\colon\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))\to\mathrm{Iso}(V\otimes\mathbb{C})\cong O(\ell)\rtimes\mathbb{C}^{\ell}.$ (4.38)
Thus it follows from (4.37) that
$\mathrm{Mono}(\overline{\mathcal{M}})=r(\pi_{1}(\overline{\mathcal{M}}\setminus\overline{\Sigma}))~{}\leq~{}r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))),$ (4.39)
i.e., the monodromy group of $\overline{\mathcal{M}}$ is a subgroup of $r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))$. We are to show in what follows that $r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))$ is isomorphic to $\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$.
Recall that
@@
is a quotient map. By eliminating a divisor, we obtain a covering map
$\varphi_{0}\colon V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})\to\mathcal{M}\setminus(D\cup\Sigma^{\prime}).$ (4.41)
Since $\mathcal{U}(\mathcal{M})$ is the universal covering space of $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$, it is also a covering space of $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$, and the covering map is given by
$\hat{\mathbf{x}}\bigr{|}_{\lambda=0}\colon\mathcal{U}(\mathcal{M})\to V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime}),\quad p\mapsto(x^{1}\bigr{|}_{\lambda=0}(p),\cdots,x^{\ell}\bigr{|}_{\lambda=0}(p)).$ (4.42)
In conclusion, we have the following commutative diagram of covering spaces:
$\mathcal{U}(\mathcal{M})$¯$\overline{\mathcal{M}}\setminus\overline{\Sigma}$$\mathcal{M}\setminus(D\cup\Sigma^{\prime})$.$V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$$\mathrm{pr}_{1}$$\overline{\pi}$$\mathrm{pr}_{2}$$\hat{\mathbf{x}}\bigr{|}_{\lambda=0}$$\varphi_{0}$
All the maps in the diagram are covering maps. From the property of universal covering spaces, we have
$\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))=\mathrm{Deck}\left(\frac{\mathcal{U}(\mathcal{M})}{\mathcal{M}\setminus(D\cup\Sigma^{\prime})}\right).$ (4.43)
Here $\mathrm{Deck}(X/Y)$ is the deck transformation group on $X$ over $Y$, where $X$ is a covering space over $Y$. Relation (4.43) implies that, every element $\gamma\in\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))$ induces a deck transformation on $\mathcal{U}(\mathcal{M})$. We have the following lemma for $r(\gamma)$.
@@
Let $\gamma\in\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))$, and let us view $\gamma$ as a deck transformation on $\mathcal{U}(\mathcal{M})$. Then $p=r(\gamma)$ is the unique deck transformation on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$ over $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$ such that
$\hat{\mathbf{x}}\bigr{|}_{\lambda=0}\circ\gamma=p\circ\hat{\mathbf{x}}\bigr{|}_{\lambda=0}.$ (4.44)
In other words, we have the following commutative diagram.
$V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$$V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$$\mathcal{M}\setminus(D\cup\Sigma^{\prime})$.$\mathcal{U}(\mathcal{M})$$\mathcal{U}(\mathcal{M})$$\exists!~{}p=r(\gamma)$$\varphi_{0}$$\varphi_{0}$$\gamma$$\hat{\mathbf{x}}\Big{|}_{\lambda=0}$$\hat{\mathbf{x}}\Big{|}_{\lambda=0}$
@@
Since $\hat{\mathbf{x}}\bigr{|}_{\lambda=0}$ is surjective, the map $p$ satisfying (4.44) is uniquely determined. Moreover, since $r(\gamma)=\gamma^{*}$ is defined as analytic continuation on $\mathcal{U}(\mathcal{M})$, $p=r(\gamma)$ satisfies (4.44). Thus in order to prove the lemma, it suffices to verify that $r(\gamma)$ is a deck transformation over $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$, i.e., $\varphi_{0}\circ r(\gamma)=\varphi_{0}$. As $\gamma$ is a deck transformation over $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$, we have
$\varphi_{0}\circ\hat{\mathbf{x}}\bigr{|}_{\lambda=0}=\varphi_{0}\circ\hat{\mathbf{x}}\Big{|}_{\lambda=0}\circ\gamma=\varphi_{0}\circ r(\gamma)\circ\hat{\mathbf{x}}\bigr{|}_{\lambda=0}.$ (4.45)
Since $\hat{\mathbf{x}}\bigr{|}_{\lambda=0}$ is surjective, it follows that
@@
Therefore, $r(\gamma)$ is a deck transformation on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$ over $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$. The lemma is proved. ∎
@@
The group $r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))$ can be represented as
$r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))=\mathrm{Iso}(V\otimes\mathbb{C})\cap\mathrm{Deck}\bigg{(}\frac{V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})}{\mathcal{M}\setminus(D\cup\Sigma^{\prime})}\bigg{)}.$ (4.46)
The right hand side consists of all the isometries on $V\otimes\mathbb{C}$ whose restrictions on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$ are deck transformations over $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$.
@@
For any $\gamma\in\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))$, $r(\gamma)$ is of course in $\mathrm{Iso}(V\otimes\mathbb{C})$. Moreover, by using Lemma 4.4, we know that $r(\gamma)$ is a deck transformation on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$ over $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$.
On the other hand, suppose $p$ be an isometry of $V\otimes\mathbb{C}$ and also be a deck transformation on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$ over $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$. By the general lifting lemma of covering spaces, we know that the deck transformation $p$ can be lifted to a deck transformation $\sigma:\mathcal{U}(\mathcal{M})\to\mathcal{U}(\mathcal{M})$, since
$p_{*}\left(\hat{\mathbf{x}}\bigr{|}_{\lambda=0}\right)_{*}\pi_{1}(\mathcal{U}(\mathcal{M}))=\{e\}.$ (4.47)
Therefore, we have the following commutative diagram:
$V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$$V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$$\mathcal{M}\setminus(D\cup\Sigma^{\prime})$.$\mathcal{U}(\mathcal{M})$$\mathcal{U}(\mathcal{M})$$p$$\varphi_{0}$$\varphi_{0}$$\sigma$$\hat{\mathbf{x}}\bigr{|}_{\lambda=0}$$\hat{\mathbf{x}}\bigr{|}_{\lambda=0}$

### v3

@@ 4.3. The Monodromy group of $\mathcal{M}(R,\omega)$
and $E^{\ell}$ be the dual space of $\mathcal{S}$. Then $v^{1},\dots,v^{\ell}$ span $\mathcal{S}$, and they also give a coordinate system on $E^{\ell}$. The transformation (4.33) yields an affine transformation on $E^{\ell}$, which we also denote by $\gamma^{*}$. This transformation on $E^{\ell}$ is an isometry w.r.t. the metric $g$, hence we obtain a representation
$r\colon\pi_{1}(M\setminus\Sigma)\to\mathrm{Iso}(E^{\ell})\cong O(\ell,\mathbb{C})\ltimes\mathbb{C}^{\ell},\quad\gamma\mapsto\gamma^{*}$ (4.35)
of the fundamental group of $M\setminus\Sigma$.
@@
$\mathrm{Mono}(M)=r(\pi_{1}(\mathcal{M}\setminus\Sigma))\leq O(\ell,\mathbb{C})\ltimes\mathbb{C}^{\ell}.$ (4.36)
In our construction, the inner product space $E^{\ell}$ can be identified with $V\otimes\mathbb{C}$. Let
$\Sigma=\left\{z\in\mathcal{M}_{D}\mid\det\left(g^{ij}(z)\right)=0\right\},$
and $\mathcal{U}(\mathcal{M}_{D})$ be the universal covering space of $\mathcal{M}_{D}\setminus\Sigma$. Then the flat coordinates of $g$ are $x^{1}\bigr|_{\lambda=0},\cdots,x^{\ell}\bigr|_{\lambda=0}$, which are functions on $\mathcal{U}(\mathcal{M}_{D})$. Moreover, $x^{1}\bigr|_{\lambda=0},\cdots,x^{\ell}\bigr|_{\lambda=0}$ are also coordinates of $V\otimes\mathbb{C}$, so the complexified Euclidean space $E^{\ell}$ is canonically isomorphic to $V\otimes\mathbb{C}$. Hence we obtain a representation
$r\colon\pi_{1}(\mathcal{M}_{D}\setminus\Sigma)\to\mathrm{Iso}(V\otimes\mathbb{C})\cong O(\ell,\mathbb{C})\ltimes\mathbb{C}^{\ell},\quad\gamma\mapsto\gamma^{*},$ (4.37)
@@
Here the image $r(\pi_{1}(\mathcal{M}_{D}\setminus\Sigma))$ is the monodromy group of $\mathcal{M}_{D}$. Now we proceed to compute the monodromy group of the generalized Frobenius manifolds $\mathcal{M}_{D}$ and $\widehat{\mathcal{M}}_{D}$. We will end this subsection by proving that
$\mathrm{Mono}(\mathcal{M}_{D})\cong\mathrm{Stab}_{W}(\omega)\ltimes\mathbb{Z}^{\ell}$
and $\mathrm{Mono}(\widehat{\mathcal{M}}_{D})$ is a subgroup of $\mathrm{Mono}(\mathcal{M}_{D})$.
Proposition 4.8.
There exists a natural embedding $\mathrm{Mono}(\widehat{\mathcal{M}}_{D})\hookrightarrow\mathrm{Mono}(\mathcal{M}_{D})$.
Proof.
@@
$\widehat{\pi}\colon\widehat{\mathcal{M}}_{D}\to\mathcal{M}_{D}$ (4.39)
is nondegerate, so from
$\det\left(g^{\alpha\beta}(t)\right)=\det\left(g^{ij}(z)\right)\det\left(\frac{\partial t^{\alpha}}{\partial z^{j}}\right)^{2}$ (4.40)
it follows that $\left(g^{\alpha\beta}(t)\right)$ is degenerate if and only if $\left(g^{ij}(z(t))\right)$ is degenerate. In other words,
$(\widehat{\pi})^{-1}(\Sigma)=\widehat{\Sigma}:=\left\{t\in\widehat{\mathcal{M}}_{D}\mid\det\bigl(g^{\alpha\beta}(t)\bigr)=0\right\},$
and the map
$\widehat{\pi}:\widehat{\mathcal{M}}_{D}\setminus\widehat{\Sigma}\to\mathcal{M}_{D}\setminus\Sigma$ (4.41)
[…]
$\mathrm{Mono}(\widehat{\mathcal{M}}_{D})\hookrightarrow\mathrm{Mono}(\mathcal{M}_{D})$ (4.42)
as subgroups of $\mathrm{Iso}(V\otimes\mathbb{C})$. The proposition is proved. ∎
Now we proceed to show that $\mathrm{Mono}(\mathcal{M}_{D})$ is isomorphic to $\mathrm{Stab}_{W}(\omega)\ltimes\mathbb{Z}^{\ell}$. Recall that
@@
is a branched covering, which induces an unbranched covering map
$\varphi_{0}\colon V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma)\to\mathcal{M}_{D}\setminus\Sigma.$ (4.44)
Since $\mathcal{U}(\mathcal{M}_{D})$ is the universal covering space of $\mathcal{M}_{D}\setminus\Sigma$, it is also the universal covering space of $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma)$. We denote the corresponding covering map by
$j_{0}\colon\mathcal{U}(\mathcal{M}_{D})\to V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma),\quad s\mapsto(x^{1}\bigr|_{\lambda=0}(s),\cdots,x^{\ell}\bigr|_{\lambda=0}(s)).$ (4.45)
We have the following commutative diagram of covering spaces:
$\mathcal{U}(\mathcal{M}_{D})$$\widehat{\mathcal{M}}_{D}\setminus\widehat{\Sigma}$$\mathcal{M}_{D}\setminus\Sigma$$V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma)$$\widehat{\pi}$$j_{0}$$\varphi_{0}$
All the maps in the diagram are covering maps.
To determine the image of $\pi_{1}(\mathcal{M}_{D}\setminus\Sigma)$ under the map $r$ that is introduced in (4.37), we are to use the isomorphism
$\pi_{1}(\mathcal{M}_{D}\setminus\Sigma)\cong\mathrm{Deck}\left(\frac{\mathcal{U}(\mathcal{M}_{D})}{\mathcal{M}_{D}\setminus\Sigma}\right),$ (4.46)
here $\mathrm{Deck}(X/Y)$ is the deck transformation group on the covering space $X$ over $Y$. We will also denote by $\gamma$ the deck transformation on $\mathcal{U}(\mathcal{M}_{D})$ induced by $\gamma\in\pi_{1}(\mathcal{M}_{D}\setminus\Sigma)$.
@@
Let $\gamma\in\pi_{1}(\mathcal{M}_{D}\setminus\Sigma)$, then $p=r(\gamma)$ is the unique deck transformation on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma)$ over $\mathcal{M}_{D}\setminus\Sigma$ such that
$j_{0}\circ\gamma=p\circ j_{0}.$ (4.47)
In other words, we have the following commutative diagram:
[…]
@@
The uniqueness follows from the fact that the covering map $j_{0}$ is surjective. Now we show that $p=r(\gamma)$ satisfies the commutative diagram. Since $r(\gamma)=\gamma^{*}$ is defined as analytic continuation on $\mathcal{U}(\mathcal{M}_{D})$, $p=r(\gamma)$ satisfies (4.47). We then only need to verify that $r(\gamma)$ is a deck transformation on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma)$ over $\mathcal{M}_{D}\setminus\Sigma$, i.e., $\varphi_{0}\circ r(\gamma)=\varphi_{0}$. Since $\gamma$ is a deck transformation over $\mathcal{M}_{D}\setminus\Sigma$, we have
$\varphi_{0}\circ j_{0}=\varphi_{0}\circ j_{0}\circ\gamma=\varphi_{0}\circ r(\gamma)\circ j_{0}.$ (4.48)
Since $j_{0}$ is surjective, it follows that
@@
The lemma is proved. ∎
@@
The monodromy group $r(\pi_{1}(\mathcal{M}_{D}\setminus\Sigma))$ can be represented as
$r(\pi_{1}(\mathcal{M}_{D}\setminus\Sigma))=\mathrm{Deck}\bigg(\frac{V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma)}{\mathcal{M}_{D}\setminus\Sigma}\bigg).$ (4.49)
In particular, every deck transformation on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma)$ over $\mathcal{M}_{D}\setminus\Sigma$ can be extended to an isometry on $V\otimes\mathbb{C}$.
@@
For any $\gamma\in\pi_{1}(\mathcal{M}_{D}\setminus\Sigma)$, we know from Lemma 4.9 that $r(\gamma)$ is a deck transformation on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma)$ over $\mathcal{M}_{D}\setminus\Sigma$.
On the other hand, suppose $p$ is a deck transformation on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma)$ over $\mathcal{M}_{D}\setminus\Sigma$. Since $\mathcal{U}(\mathcal{M}_{D})$ is their universal covering space, we know that the deck transformation $p$ can be lifted to a deck transformation $\sigma:\mathcal{U}(\mathcal{M}_{D})\to\mathcal{U}(\mathcal{M}_{D})$. Therefore, we have the following commutative diagram:
[…]